[Hallc_running] Hall C status 1010 Saturday

Jay Benesch benesch at jlab.org
Sat Mar 21 10:20:46 EDT 2020

I know only what's in HCLOG since I haven't talked to anyone and I don't 
have the training needed to enter the hall under present laser 
conditions.  See log entries linked below.  I imagine that laser 
alignment will take the rest of day shift and target alignment will 
resume thereafter.  I said at the PD meeting that AFAIK techs are still 
scheduled for close-up Sunday morning with RadCon support.  When I said 
installation seemed to be ahead of schedule Dave Hamlette frowned so I 
didn't bring up the idea of advancing that work.

Hall C beam line vacuum is mediocre on diagnostic girder and tripped 
thereafter.  Turbo pumps should still be on the upstream and downstream 
beam lines since Joshua Thomason isn't laser trained and so can't remove 
them today.  This needs to get done before PSS sweep by Ops since the 
equipment can only be removed in restricted access.

There was a PD meeting at 0900 this morning, Bluejeans 569 476 853 . 
Stuart was on the line.  Eight of the ten DOE labs are shut down via 
state governor orders, all but JLab and ORNL.  DOE is having near-daily 
discussions with lab directors per Stuart.  I got the impression that he 
thought Gov. Northam would pull the plug before DOE did.  Stuart asked 
that the running experiments take whatever calibration runs needed 
sooner rather than later and asked situations.  Halls A and D are in 
good shape.  Stepan said that B just switched to a new RTP and was 
counting on calibration late in the run [April 23-25 per schedule].  I 
said that one of a pair of experiments approved in 2006 with the He3 
target had just completed and that the second was in preparation and 
needed at least six weeks of data.

Ken Baggett spoke about the meeting Steve Suhring has called for 1330 
Monday (Bluejeans 569 476 853 ) would cover Steve's agenda and 
differences to shutdown checklists prepared for hurricanes.  This 
shutdown could be months per Ken.  I asked Stuart whether there was any 
long term planning given that the situation could last 18-24 months.  He 
said he was sure DOE folks were considering that.

HCLOG entries

Target ladder alignment progress

Cell "Austin" PC Position Check and Laser Alignment

Partial progress of laser work

Steve Suhring's email re Monday 1330

We will meet Monday at 1330 on-line.

The following link will take you to the lab tool for virtual meetings. 
As you log on, be sure to mute your microphone to avoid an echo-like 
audio loop.



1. Changes to the plan since our last meeting

2. Possible outcomes for Summer work and the variables contributing to 
the SAD Planning

3. Areas needing attention leading to the 2020SAD

4. Other?

More information about the Hallc_running mailing list